# Anatomy of singlet-doublet dark matter relic: annihilation, co-annihilation, co-scattering, and freeze-in

### v1

@@ 5 Displaced vertex signatures
Because of gauge interactions, the doublet components, $\psi^{\pm}$, and $\chi_{0}$ can be copiously produced in the colliders. Once these particles are produced, they will decay to the DM and charged lepton via the off-shell gauge bosons in the small mass splitting range after traveling some finite distance. The decay length is $\propto(\sin^{2}\theta\Delta{M}^{5})^{-1}$. Thus, one can measure the displaced vertex signatures in present and future collider experiments such as LHC and MATHUSLA. The inclusion of co-scattering can enhance the reach of these experiments in searching for these dark sector particles. We have calculated the decay lengths of such particles considering mass splitting ($\Delta M$) in the range from 1 to 12 GeV and shown in Fig. 10 with varying DM mass. We considered three ranges of mixing angles: $10^{-2}\leq\sin\theta\leq 2\times 10^{-2}$, $10^{-4}\leq\sin\theta\leq 2\times 10^{-4}$, and $10^{-5}\leq\sin\theta\leq 2\times 10^{-5}$. The three patches correspond to these mixing angles in the bottom to top order, respectively. The color code represents the parameter $\rho_{2}$. The figure depicts that the inclusion of the co-scattering enhances the lifetime of the dark partners and hence reaches the decay length of LHC and MATHUSLA.
Figure 10: Decay length of the doublet fermion as a function of DM mass with $\rho_{2}$ in the color code. The mixing angles are varied in the range of $10^{-2}\leq\sin\theta\leq 2\times 10^{-2}$, $10^{-4}\leq\sin\theta\leq 2\times 10^{-4}$, and $10^{-5}\leq\sin\theta\leq 2\times 10^{-5}$ and the three patches correspond to these mixing angles in the bottom to top order, respectively.
@@ 6 Conclusions
In this work, we explored the full parameter space of singlet-doublet vector-like fermionic dark matter for the first time, accounting for annihilation, co-annihilation, co-scattering, and freeze-in processes. In Fig. 11, we show the parameter space in the plane of mixing angle versus DM mass, where the black dashed line separates the thermal and non-thermal regimes at $\Delta M$ = 1 GeV. For $\Delta M<$ 1 GeV, this line shifts downward, and for $\Delta M>$ 1 GeV, this line moves upward. In the upper part of the line, the DM can attain thermal equilibrium, and the relic can be obtained via freeze-out mechanism. On the other hand, below this black dashed line, the DM can never reach equilibrium, and the relic can be obtained via the freeze-in mechanism only. The grey-shaded region represents the parameter space where the annihilation and co-annihilation processes solely decide the relic. When the mixing angle drops further, the co-scattering processes become the most dominant one, and they decide the DM relic in this regime. This is shown with the green-shaded region.
As $\sin\theta$ decreases further, the DM decouples from equilibrium earlier, freezing out with a larger abundance. This over-abundant region is depicted as the cyan-shaded area. When the mixing angle drops below $<2\times 10^{-7}$, thermal production of DM becomes inefficient. Below the black dashed line, the doublet decays into the singlet DM, which subsequently freezes into the correct relic abundance. This freeze-in regime is shown in the blue-shaded region. However, for $\sin\theta<\mathcal{O}(10^{-13})$, the doublet components fail to produce enough singlet DM, resulting in an under-produced region shown in white.
@@ 7 Acknowledgments
Appendix A Evolution of abundances for point A
[…]
Appendix B Decay rates
@@
$\Gamma_{a\rightarrow bc}$ $=$ $\frac{\sqrt{\Lambda(M_{a}^{2},M_{b}^{2},M_{c}^{2})}}{2M_{a}}\frac{1}{32\pi^{2}M_{a}^{2}}4\pi\left|\mathcal{M}_{a\rightarrow bc}\right|^{2},$ (21)
${\rm where,~{}}\Lambda(x,y,z)$ $=$ $x^{2}+y^{2}+z^{2}-2xy-2yz-2xz$ (22)
$\left|\mathcal{M}_{\psi^{0}\rightarrow\chi H}\right|^{2}$ $=$ $\frac{(M_{\chi_{0}}-M_{\chi_{1}})^{2}\sin^{2}2\theta}{4v^{2}}\left(1-2\sin^{2}\theta\right)^{2}\left(M_{\chi_{1}}^{2}+2M_{\chi_{0}}M_{\chi_{1}}+M_{\chi_{0}}^{2}-M_{h}^{2}\right),$ (23)
$\left|\mathcal{M}_{\psi^{0}\rightarrow\chi Z}\right|^{2}$ $=$ $G_{F}\left(1-\sin^{2}\theta\right)\sin^{2}\theta(M^{4}_{\chi_{1}}-2M^{2}_{\chi_{0}}M^{2}_{\chi_{1}}+M^{2}_{\chi_{1}}M^{2}_{Z}-6M_{\chi_{0}}M_{\chi_{1}}M^{2}_{Z}+M^{4}_{\chi_{0}}$ (24)
@@
$\left|\mathcal{M}_{\psi^{-}\rightarrow\chi W}\right|^{2}$ $=$ $\frac{G_{F}}{2}\sin^{2}\theta(M^{4}_{\chi_{1}}-2M^{2}_{\psi^{-}}M^{2}_{\chi_{1}}+M^{2}_{\chi_{1}}M^{2}_{W}-6M_{\psi^{-}}M_{\chi_{1}}M^{2}_{W}+M^{4}_{\psi^{-}}$ (25)
@@
$\Gamma_{\psi\rightarrow\chi l\nu_{l}}=\frac{2G_{F}^{2}}{15\pi^{3}}(\Delta M)^{5}$ (26)

### v2

@@ 5 Displaced vertex signatures
Figure 13: Decay length of the doublet fermion as a function of DM mass with $\rho_{2}$ in the color code. The mixing angles are varied in the ranges: $10^{-2}\leq\sin\theta\leq 2\times 10^{-2}$ (bottom), $10^{-4}\leq\sin\theta\leq 2\times 10^{-4}$ (middle), and $10^{-5}\leq\sin\theta\leq 2\times 10^{-5}$ (top).
Because of gauge interactions, the doublet components, $\psi^{\pm}$, and $\chi_{0}$ can be copiously produced in the colliders. Once these particles are produced, they will decay to the DM and charged lepton via the off-shell gauge bosons in the small mass splitting range after traveling some finite distance. The decay length is $\propto(\sin^{2}\theta\Delta{M}^{5})^{-1}$. Thus, one can measure the displaced vertex signatures in present and future collider experiments such as LHC and MATHUSLA. The inclusion of co-scattering can enhance the reach of these experiments in searching for these dark sector particles. We have calculated the decay lengths of such particles considering mass splitting ($\Delta M$) in the range from 1 to 12 GeV and shown in Fig. 13 with varying DM mass. We considered three ranges of mixing angles: $10^{-2}\leq\sin\theta\leq 2\times 10^{-2}$ (top), $10^{-4}\leq\sin\theta\leq 2\times 10^{-4}$ (middle), and $10^{-5}\leq\sin\theta\leq 2\times 10^{-5}$ (bottom). The color code represents the parameter $\rho_{2}$. The figure depicts that the inclusion of the co-scattering enhances the lifetime of the dark partners and hence reaches within the sensitivities of LHC and MATHUSLA.
@@ 6 Conclusions
In this work, we explored the full parameter space of singlet-doublet vector-like fermionic dark matter for the first time, accounting for annihilation, co-annihilation, co-scattering, and decay-inverse decay processes for both freeze-out and freeze-in production of DM along with the SuperWIMP mechanism. In Fig. 14, we show the parameter space in the plane of mixing angle versus DM mass, where the black dashed line separates the thermal and non-thermal regimes at $\Delta M$ = 1 GeV. For $\Delta M<$ 1 GeV, this line shifts downward, and for $\Delta M>$ 1 GeV, this line moves upward. In the region above the black dashed line, the DM can attain thermal equilibrium, and the relic can be obtained via freeze-out mechanism. On the other hand, below this black dashed line, the DM can never reaches equilibrium, and the relic can be obtained via the non-thermal processes only. The gray-shaded region represents the parameter space where the annihilation and co-annihilation processes solely decide the relic. When the mixing angle is reduced further, the co-scattering processes become the most dominant one, and they decide the DM relic in this regime. This is shown with the green-shaded region. As $\sin\theta$ decreases further, the DM decouples from equilibrium earlier, freezing out with a larger abundance. This over-abundant region is depicted as the cyan-shaded area. When the mixing angle drops below $2\times 10^{-7}$, the DM never reaches equilibrium. Below the black dashed line, $\sin\theta$ up to $\mathcal{O}(10^{-11})$, DM relic gets over populated by the freeze-in channels. In the range, $\mathcal{O}(10^{-11})\gtrsim\sin\theta\gtrsim\mathcal{O}(10^{-13})$, the DM relic can be achieved via the freeze-in mechanism. For further smaller $\sin\theta$, typically $\mathcal{O}(10^{-14})$, the DM relic can be obtained through the SuperWIMP mechanism.
@@ 7 Acknowledgments
Appendix A Evolution of abundances for point $A$ in Table 1
Here we would like to analyze the dependency of the relic for point $A$ in Table 1 on the decay and inverse decay terms as given in Eq. (3.8) by switching off the co-scattering term. In Fig. 15, we have shown $Y_{1},Y_{2}$ as a function of temperature. The parameters are mentioned in the figure inset.
Figure 15: Left: Evolution of abundances for the point $A$ in Table 1 without considering decay. As the $\sin\theta$ is small, once the co-annihilation processes decouple, the singlet decouples early with a larger abundance while the doublet remains in equilibrium for a longer epoch resulting in a smaller abundance. Right: Evolution of abundances for the point $A$ in Table 1 including decay. Here in presence of decay and inverse decay, the equilibration is maintained for longer duration. This results in correct relic density of the DM.
Appendix B Feynman Diagrams of the processes involved in relic density
B.1 Annihilation of dark matter
The sector 1 particle (i.e. the DM) annihilate to the SM particles and the relevant processes are listed in Fig. 16.
Figure 16: DM annihilating to the SM particles through 1100 processes.
B.2 Annihilation and co-annihilation among the sector 2 particles
The sector 2 particles (i.e. the doublet components) annihilate and co-annihilate to the SM particles and the relevant processes are listed in Fig. 17.
Figure 17: Annihilation and co-annihilation of the doublet components to the SM particles through 2200 processes.
B.3 Co-annihilation processes
The sector 1 particle co-annihilate with the sector 2 particles to the SM particles. The relevant processes are listed in Fig. 18 and Fig. 19.
Figure 18: DM co-annihilating with the neutral component of the doublet through 1200 processes.
Figure 19: DM co-annihilating with the charged component of the doublet through 1200 processes.
B.4 Co-scattering processes
The sector 1 particle scatters with the SM particles to the sector 2 particles and SM particles known as the co-scattering processes. The relevant processes are listed in Fig. 20, and Fig. 21.
Figure 20: DM co-scattering with the SM bath particles to the neutral component of the doublet state through 1020 processes.
Figure 21: DM co-scattering with the SM bath particles to the charged component of the doublet state through 1020 processes.
Appendix C Decay rates
The kinematically allowed two and three body decay modes of the charged and neutral component of the doublet are shown in Fig. 22.
@@ Appendix C Decay rates
$\Gamma_{a\rightarrow bc}$ $=$ $\frac{\sqrt{\Lambda(M_{a}^{2},M_{b}^{2},M_{c}^{2})}}{2M_{a}}\frac{1}{32\pi^{2}M_{a}^{2}}4\pi\left|\mathcal{M}_{a\rightarrow bc}\right|^{2},$ (C.1)
${\rm where,\leavevmode\nobreak\ }\Lambda(x,y,z)$ $=$ $x^{2}+y^{2}+z^{2}-2xy-2yz-2xz$ (C.2)
$\left|\mathcal{M}_{\psi^{0}\rightarrow\chi H}\right|^{2}$ $=$ $\frac{(M_{\chi_{0}}-M_{\chi_{1}})^{2}\sin^{2}2\theta}{4v^{2}}\left(1-2\sin^{2}\theta\right)^{2}\left(M_{\chi_{1}}^{2}+2M_{\chi_{0}}M_{\chi_{1}}+M_{\chi_{0}}^{2}-M_{h}^{2}\right),$ (C.3)
$\left|\mathcal{M}_{\psi^{0}\rightarrow\chi Z}\right|^{2}$ $=$ $G_{F}\left(1-\sin^{2}\theta\right)\sin^{2}\theta(M^{4}_{\chi_{1}}-2M^{2}_{\chi_{0}}M^{2}_{\chi_{1}}+M^{2}_{\chi_{1}}M^{2}_{Z}-6M_{\chi_{0}}M_{\chi_{1}}M^{2}_{Z}+M^{4}_{\chi_{0}}$ (C.4)
@@
$\left|\mathcal{M}_{\psi^{-}\rightarrow\chi W}\right|^{2}$ $=$ $\frac{G_{F}}{2}\sin^{2}\theta(M^{4}_{\chi_{1}}-2M^{2}_{\psi^{-}}M^{2}_{\chi_{1}}+M^{2}_{\chi_{1}}M^{2}_{W}-6M_{\psi^{-}}M_{\chi_{1}}M^{2}_{W}+M^{4}_{\psi^{-}}$ (C.5)
@@
$\Gamma_{\psi\rightarrow\chi l\nu_{l}}=\frac{2G_{F}^{2}}{15\pi^{3}}(\Delta M)^{5}$ (C.6)